mathslover (mathslover):

Find the minimum number of roots of the equation : \[\large{p(x)= x^10 + 24x^8 -35x^7+5x^6+25x^5+3x^3+4=0}\]

mathslover (mathslover):

Which are imaginary ...

OpenStudy (maheshmeghwal9):

Is the first variable\[\LARGE \color{red}{x^{10}}\]?

mathslover (mathslover):

yes sorry i forgot to put that in brackets

mathslover (mathslover):

\[\large{\color{violetred}{p(x)=x^{10}+24x^8−35x^7+5x^6+25x^5+3x^3+4=0}}\]

OpenStudy (maheshmeghwal9):

I have done questions like this but I have forgotten that:( Can u remind me what concept this question is using?

OpenStudy (anonymous):

the mane theory of algebra states that the number of roots of a polinomial is equal to it's degree. In this case degree is 10, so 10 roots

mathslover (mathslover):

@myko it is asked for imaginary roots

mathslover (mathslover):

i just only know about replacing x by (-x) in the equation p(x)

mathslover (mathslover):

it is written that if x is replaced by (-x) then the max. number of negative roots for the given equation is given by number of changes in the sign of the equation for -ve roots

mathslover (mathslover):

@maheshmeghwal9 i only know this theory

mathslover (mathslover):

can u give me any link that where i can learn this ..

OpenStudy (anonymous):

http://www.purplemath.com/modules/drofsign.htm

mathslover (mathslover):

from that also i cant get the answer...

OpenStudy (maheshmeghwal9):

If u don't mind I will post the answer after 2 or 2.5c hrs:)

OpenStudy (maheshmeghwal9):

100%

OpenStudy (maheshmeghwal9):

& i hope u gt the answer before i give u:)

mathslover (mathslover):

ok .. no problem hope for best..

mathslover (mathslover):

i got the solution is it correct can any 1 help me in this : \[\large{p(x)=x^{10}+24x^8-35x^7+5x^6+25x^5+3x^3+4=0}\]

mathslover (mathslover):

wait..

mathslover (mathslover):

is this answer correct ?

mathslover (mathslover):

@maheshmeghwal9 and @ganeshie8 , @shubhamsrg ...

OpenStudy (shubhamsrg):

nops..the actual ans is it has 10 imaginary roots..(source : wolframalpha) hmmn and the theorem you used works only when you know that roots are real,,just need to know there sign,,

mathslover (mathslover):

can u give me the link of that question in wolfram alpha ? @shubhamsrg

mathslover (mathslover):

i had calculated min. possible imaginary roots..

mathslover (mathslover):

Minimum number of imaginary roots : 10-4 = 6

OpenStudy (shubhamsrg):

what do you mean by min no. of possible roots?

OpenStudy (shubhamsrg):

ohh..i understand now.. hmm,,yes,,your sol might be correct @mathslover

OpenStudy (maheshmeghwal9):

@mathslover I have found the way of finding the minimum number of possible imaginary roots of this polynomial. But I m having problem in taking second, third derivatives.

mathslover (mathslover):

@maheshmeghwal9 there is no need of deriviatives here ... http://assets.openstudy.com/updates/attachments/4fddb20ee4b0f2662fd2ff1a-mathslover-1339932263149-roots.png

OpenStudy (maheshmeghwal9):

no I m doing this question by STURM's Theorem so wanna take derivatives. @mathslover

mathslover (mathslover):

but why r u making this more complicated .... well i know that may be important for IIT ...

mathslover (mathslover):

where r u having the problems in deriviatives post them

OpenStudy (maheshmeghwal9):

k! thats ur wish. but sturm's theorem is simple dude. k bye:)

mathslover (mathslover):

bbye ...

OpenStudy (maheshmeghwal9):

ur answer seems to be right by sturm's theorem too:)

mathslover (mathslover):

:)

mathslover (mathslover):

thanks

OpenStudy (maheshmeghwal9):

yw:)